OpenStudy (anonymous):

Put 5x^2+7x-3 in the form a(x+h)+k My answer (x+7/10)^2-109/20. I need help finding the a value.

OpenStudy (anonymous):

\(\large 5x^2+7x-3 \) \(\large =5[x^2+\frac{7}{5}x]-3 \) \(\large =5[x^2+\frac{7}{5}x+\frac{49}{100}-\frac{49}{100}]-3 \) \(\large =5[(x^2+\frac{7}{5}x+\frac{49}{100})-\frac{49}{100}]-3 \) \(\large =5[(x+\frac{7}{10})^2-\frac{49}{100}]-3 \) \(\large =5(x+\frac{7}{10})^2-5\cdot\frac{49}{100}-3 \) \(\large =5(x+\frac{7}{10})^2-\frac{49}{20}-\frac{60}{20} \) \(\large =5(x+\frac{7}{10})^2-\frac{109}{20} \) so your "a" value should be 5